# Dissecting Logical Reasoning in LLMs: A Fine-Grained Evaluation and Supervision Study

### v1

@@ Abstract
Abstract: Logical reasoning is a core capability for many applications of large language models (LLMs), yet existing benchmarks often rely solely on final-answer accuracy, failing to capture the quality and structure of the reasoning process. We propose FineLogic, a fine-grained evaluation framework that assesses logical reasoning across three dimensions: overall benchmark accuracy, stepwise soundness, and representation-level alignment. In addition, to better understand how reasoning capabilities emerge, we conduct a comprehensive study on the effects of supervision format during fine-tuning. We construct four supervision styles (one natural language and three symbolic variants) and train LLMs under each. Our findings reveal that natural language supervision yields strong generalization even on out-of-distribution and long-context tasks, while symbolic reasoning styles promote more structurally sound and atomic inference chains. Further, our representation-level probing shows that fine-tuning primarily improves reasoning behaviors through step-by-step generation, rather than enhancing shortcut prediction or internalized correctness. Together, our framework and analysis provide a more rigorous and interpretable lens for evaluating and improving logical reasoning in LLMs. The code is available at https://github.com/YujunZhou/Logical.
@@ 1 Introduction
Large language models (LLMs) are rapidly emerging as transformative tools across a wide array of applications (Achiam et al., 2023; Guo et al., 2024b; Thirunavukarasu et al., 2023; Nam et al., 2024). Among these, reasoning serves as a core capability underpinning tasks such as problem-solving (Lu et al., 2023), scientific question answering (Guo et al., 2024a), and code analysis (Nam et al., 2024). Consequently, a growing body of research has sought to evaluate and enhance the reasoning abilities of LLMs from multiple perspectives (Wei et al., 2022; Guo et al., 2025, 2024a). Within this broader landscape, logical reasoning stands out as a particularly challenging and intellectually demanding domain (Saparov and He, 2022). It requires a synthesis of natural language understanding, formal logical interpretation, and multi-step inferential processing (Patel et al., 2024; Saparov et al., 2023; Morishita et al., 2024).
@@
Despite growing interest in the logical reasoning capabilities of LLMs, most existing benchmarks focus narrowly on whether a model produces the correct final answer (Patel et al., 2024; Parmar et al., 2024; Han et al., 2022). This binary evaluation, typically assessing only the correctness of a “True” or “False” output, can be misleading, as it fails to determine whether the model arrived at the answer through valid multi-step reasoning (Saparov and He, 2022). Consequently, correct answers may reflect guesswork rather than genuine logical inference. We are thus motivated to address RQ1: How to rigorously evaluate LLMs’ step-by-step correctness in logical reasoning tasks, beyond the binary evaluation of the final answer?
@@
In parallel with benchmarking efforts, numerous methods have been proposed to enhance the multi-step logical reasoning abilities of LLMs. While many leverage inference-time strategies (Wang et al., 2025), in-context learning (Creswell et al., 2022; Xu et al., 2024), or external logical verifiers (Pan et al., 2023) to guide the model toward more rigorous reasoning, some recent studies explored supervised fine-tuning (SFT) as a more direct approach to enhancing logical reasoning (Morishita et al., 2024; Feng et al., 2023). For example, Morishita et al. (2024) proposes a synthetic logic corpus designed to offer broad and systematic coverage of logical knowledge. However, it remains unclear for this important question, RQ2: What style of training data, natural language or formal logical symbols, better facilitates the learning of multi-step logical reasoning through SFT? Addressing this research question is important for understanding how to most effectively instill logical reasoning capabilities in LLMs.
@@
To address RQ1, we propose FineLogic, a new evaluation framework designed to more fine-grainedly assess the logical reasoning capabilities of LLMs. Specifically, our framework evaluates models along three complementary dimensions: (1) Overall benchmark accuracy: This metric captures a model’s ability to perform multi-step logical reasoning and its generalizability across problems from diverse domains. (2) Stepwise Soundness: Inspired by Saparov and He (2022), we assess the quality of each intermediate reasoning step using three criteria—validity (whether the step is logically valid), relevance (whether its conclusion is used in later steps), and atomicity (whether it applies a single, minimal inference rule). These metrics aim to evaluate the model’s ability to generate human-interpretable and logically coherent reasoning chains. (3) Representation-level probing (Ye et al., 2024): By applying probing techniques to LLM hidden representations, this evaluation provides insight into whether the model’s understanding of logical structure is merely surface-level or embedded in its internal state.
@@
  • We propose FineLogic, a unified and rigorous evaluation framework for assessing LLMs’ logical reasoning, moving beyond final-answer accuracy to evaluate the quality, interpretability, and coherence of their solutions.
@@
  • Through systematic analysis of models trained with different supervision styles, we identify key trade-offs between generalization and structural reasoning quality. These findings provide concrete insights into the design and selection of effective training data for supervised logical reasoning.
@@ 2 Related Works
Logical Reasoning Benchmarks. Numerous benchmarks have been proposed to evaluate the logical reasoning abilities of LLMs. Liu et al. (2023); Luo et al. (2023); Havrilla et al. (2024) mix logical and commonsense reasoning, making it hard to isolate logical competence. Others assess multi-step reasoning but rely only on final-answer accuracy (Parmar et al., 2024; Han et al., 2022; Tafjord et al., 2020; Mondorf and Plank, 2024). While ProntoQA (Saparov and He, 2022; Saparov et al., 2023) introduces stepwise evaluation, it uses short problems and focuses only on step correctness. In contrast, our FineLogic framework provides a more rigorous and comprehensive assessment across sample-level correctness, step-level reasoning quality, and internal representation alignment.
Logical Reasoning Enhancement. Several studies have aimed to improve LLMs’ performance on logical reasoning tasks. Some approaches rely on translating inputs into formal logic and using programmable verifiers to solve problems (Olausson et al., 2023; Pan et al., 2023; Yang et al., 2023; Ryu et al., 2024), which bypasses the model’s own reasoning process. Others use in-context learning or inference-time strategies to guide output without fundamentally enhancing reasoning ability (Creswell et al., 2022; Wang et al., 2025; Xu et al., 2024; Sun et al., 2023; Toroghi et al., 2024). While a few works have explored fine-tuning or reinforcement learning to strengthen logical reasoning (Feng et al., 2023; Morishita et al., 2023, 2024; Xie et al., 2025; Yang et al., 2022; Xie et al., 2024), they have not examined which types of supervision are most effective for teaching LLMs to reason. In this work, we focus specifically on this open question.
@@ 3.1 Overall Benchmark Accuracy
Similar to most benchmarks, our overall benchmark accuracy focuses on final-answer correctness. While coarse-grained, it offers a quick and effective way to assess a model’s overall reasoning ability and cross-domain generalization. We evaluate on four challenging multi-step reasoning benchmarks: FLD (Morishita et al., 2024), FOLIO (Han et al., 2022), Multi-LogiEval (Patel et al., 2024), and ProntoQA (Saparov and He, 2022). For FLD, we generate 50 samples per step (0–19) and 100 UNKNOWN cases. For FOLIO, the full test set is used. For Multi-LogiEval, we select first-order and propositional problems with depths 4–5. For ProntoQA, we follow Pan et al. (2023) and evaluate on the 500 hardest 5-hop samples. Dataset statistics are shown in Table 1, with details in Appendix A.1.
@@
FLD (Morishita et al., 2024) 1100 {T, F, Unknown}
FOLIO (Han et al., 2022) 203 {T, F, Unknown}
Multi-Logical (Patel et al., 2024) 390 {T, F}
Pronto-QA (Saparov and He, 2022) 500 {T, F}
@@ 3.2 Stepwise Soundness
To assess these criteria, we extract the premises and conclusion of each step. We use GPT-4.1-mini to evaluate validity and atomicity. Manual verification on 200 annotated steps shows that GPT-4.1-mini achieves over 98% accuracy on both metrics. For relevance, we determine whether the conclusion of step $i$ (e.g., $\texttt{int}\leavevmode\nobreak\ j$) is referenced in any subsequent step $k>i$.
We then compute the proportion of samples in which all steps are valid, relevant, and atomic, providing a sample-level measure of reasoning integrity. Full prompt templates are provided in Figures 13 and 14.
@@ 3.3 Representation-level Probing
Correctness Spanning Steps (CSS): Identifies the earliest step after which the model consistently predicts the correct label. The spanning length is the number of remaining steps from that point to the end. Higher accuracy indicates earlier internalization of the correct answer.
@@
Our evaluation builds on two prior lines of work—stepwise reasoning evaluation (Saparov and He, 2022) and representation-level probing (Ye et al., 2024)—but introduces key extensions tailored to logical reasoning.
Stepwise Soundness Evaluation.
Saparov and He (2022) evaluate reasoning steps using three criteria: validity (logical entailment), utility (contribution to the final proof), and atomicity (single rule application per step). Since utility depends on gold proof annotations and is often impractical, we propose a more accessible alternative: relevance, which checks whether a step’s conclusion is used in any subsequent inference. Moreover, prior work focuses on individual steps, while we extend this to the solution level by introducing all-steps validity, relevance, and atomicity—sample-level metrics that reflect whether a full reasoning chain is logically sound and interpretable.
Representation-Level Probing.
Ye et al. (2024) use probing to assess internal reasoning in math problems. We adapt this method to logical reasoning and introduce a new metric: Correctness Spanning Steps (CSS), which identifies the earliest point after which the model consistently predicts the correct label. CSS approximates the model’s internal reasoning depth by measuring how early it stabilizes on the correct answer.
@@ 5.1 Experimental Setup
We conduct all SFT experiments on two models: LLaMA-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, both fully fine-tuned for 3 epochs with a learning rate of $1\times 10^{-6}$.
Our baselines include four models: LLaMA-3.1, Qwen-2.5, GPT-4o, and DeepSeek R1. Fine-tuning-based methods use only LLaMA and Qwen as base models. Due to computational constraints, representation-level probing is conducted only on LLaMA, Qwen, and their SFT variants. Stepwise evaluation requires strict output formatting and enforces explicit step-by-step generation.
@@
  • Few-Shot Learning (Brown et al., 2020)
  • LOGIPT (Creswell et al., 2022) Utilizing “Pyke" and 4-turn prompting to generate semantically rich fine-tune data for downstream tasks.
  • Selection-Inference (Creswell et al., 2022) N-round selection of facts and intermediate inferences, followed by reasoning to derive new intermediate inferences.
  • SymbCoT (Xu et al., 2024) integrates symbolic expressions and first-order logic inference rules into CoT prompting and decomposing problems via a ’plan-then-solve’ pipeline.
  • LogicLM (Pan et al., 2023) introduces a framework that integrates LLMs with symbolic solvers through a three-stage process (problem formulation, symbolic reasoning, result interpretation), incorporating a self-refinement mechanism that utilizes solver feedback.
@@
Direct 31.7 54.7 40.5 64.6
CoT 29.3 50.7 44.6 63.8
Few-shot 41.0 46.5 59.4 48.9
Logic-LM 38.3 52.5 44.4 77.6
SymbCoT 38.1 58.8 46.3 78.8
Sel-Inf 48.5 47.5 55.2 64.2
LogiPT 53.3 61.7 57.9 76.4
SFT-NL 67.5 57.1 71.3 99.6
SFT-Symb-Struct 63.2 56.2 59.7 99.8
SFT-Symb-Filter 66.7 54.7 50.8 91.0
Llama-3.1-8B-Instruct SFT-Symb-Direct 52.8 48.3 53.9 98.8
@@ 5.2 Results
We conducted experiments for analyzing the performance of four models combined with various prompting and fine-tuning settings under the FineLogic Evaluation Framework.
@@ 5.2.1 Results on Overall Benchmark Accuracy
As shown in Table 1, we report the overall benchmark accuracy across four datasets, as well as the step-wise accuracy on the FLD benchmark, stratified by reasoning depth (Figure 3. Our analysis yields several key observations:
@@ 5.2.2 Results on Stepwise Soundness
The All Valid metric measures the proportion of samples in which every generated reasoning step is logically valid. This is a stringent indicator of a model’s grasp of formal reasoning rules. We observe that models trained with the SFT-NL and SFT-Symb-Struct settings achieve particularly high All Valid scores—substantially outperforming even GPT-4o and DeepSeek-R1. Notably, the Llama model fine-tuned under SFT-NL achieves an All Valid rate of 40.9
@@ 5.2.3 Results on Representation-level Probing
Table 4 presents results from our representation-level probing analysis, which aims to assess whether the models have internally acquired key reasoning abilities.
Regarding Correctness Spanning Steps (CSS), which assesses how early the model predicts the final answer, most SFT methods show little improvement. Only SFT-Symb-Filter on Llama yields a modest gain, suggesting SFT primarily guides step-by-step generation rather than enhancing early "shortcut" predictions.
For the Redundant Fact Identification metric, most SFT settings show noticeable gains. Interestingly, SFT-Symb-Direct consistently achieves the highest performance. We hypothesize that this setting, which omits the explicit logic translation phase, forces the model to implicitly learn both symbolic interpretation and reasoning. In doing so, it may develop a stronger understanding of the logical roles and dependencies among the facts, thus improving its ability to distinguish between relevant and redundant conditions.
In the Next-Step Derivability task, SFT consistently benefits Llama, particularly SFT-NL (likely due to natural language’s accessibility). In contrast, SFT shows minimal impact on Qwen, possibly because its base model is already proficient in step-tracking.

### v2

@@ Abstract
Logical reasoning is a core capability for large language models (LLMs), yet existing benchmarks that rely solely on final-answer accuracy fail to capture the quality of the reasoning process. To address this, we introduce FineLogic, a fine-grained evaluation framework that assesses logical reasoning across three dimensions: overall accuracy, stepwise soundness, and representation-level probing. Leveraging this framework, we conduct a comprehensive study on how different supervision formats in fine-tuning shape reasoning abilities. We fine-tune LLMs on four supervision styles—one in natural language and three symbolic variants—and find a key trade-off: natural language supervision excels at generalization to out-of-distribution and long-chain problems, whereas symbolic supervision is superior at instilling structurally sound, atomic reasoning steps. Furthermore, our probing analysis indicates that fine-tuning primarily refines the model’s step-by-step generation process, rather than improving its ability to converge on an answer early. Together, our framework and analysis provide a more rigorous lens for evaluating and improving logical reasoning in LLMs. The code is available at https://github.com/YujunZhou/FineLogic.
Dissecting Logical Reasoning in LLMs: A Fine-Grained Evaluation and Supervision Study
Yujun Zhou1,∗, Jiayi Ye2,∗, Zipeng Ling3,∗, Yufei Han4, Yue Huang1, Haomin Zhuang1 Zhenwen Liang1 Kehan Guo1, Taicheng Guo1, Xiangqi Wang1, Xiangliang Zhang1 1University of Notre Dame 2MBZUAI 3University of Pennsylvania 4INRIA Equal Contribution {yzhou25,xzhang33}@nd.edu
@@ 1 Introduction
Large language models (LLMs) are rapidly emerging as transformative tools across a wide array of applications Achiam et al. (2023); Guo et al. (2024b); Thirunavukarasu et al. (2023); Nam et al. (2024); Huang et al. (2024, 2025); Zhou et al. (2024a). Among these, reasoning serves as a core capability underpinning tasks such as problem-solving Lu et al. (2023), scientific question answering Guo et al. (2024a); Zhou et al. (2024b), and code analysis Nam et al. (2024). Consequently, a growing body of research has sought to evaluate and enhance the reasoning abilities of LLMs from multiple perspectives Wei et al. (2022); Guo et al. (2025, 2024a); Liang et al. (2024); Wang et al. (2025b); Zhou et al. (2025); Zhao et al. (2025). Within this broader landscape, logical reasoning stands out as a particularly challenging and intellectually demanding domain Saparov and He (2022). It requires a synthesis of natural language understanding, formal logical interpretation, and multi-step inferential processing Patel et al. (2024); Saparov et al. (2023); Morishita et al. (2024).
@@
Despite growing interest in the logical reasoning capabilities of LLMs, most existing benchmarks focus narrowly on whether a model produces the correct final answer Patel et al. (2024); Parmar et al. (2024); Han et al. (2022). This binary evaluation, typically assessing only the correctness of a “True” or “False” output, can be misleading, as it fails to determine whether the model arrived at the answer through valid multi-step reasoning Saparov and He (2022). Consequently, correct answers may reflect guesswork rather than genuine logical inference. We are thus motivated to address RQ1: How to rigorously evaluate LLMs’ step-by-step correctness in logical reasoning tasks, beyond the binary evaluation of the final answer?
@@
In parallel with benchmarking efforts, numerous methods have been proposed to enhance the multi-step logical reasoning abilities of LLMs. While many leverage inference-time strategies Wang et al. (2025a), in-context learning Creswell et al. (2022); Xu et al. (2024), or external logical verifiers Pan et al. (2023) to guide the model toward more rigorous reasoning, some recent studies explored supervised fine-tuning (SFT) as a more direct approach to enhancing logical reasoning Morishita et al. (2024); Feng et al. (2023). For example, Morishita et al. (2024) proposes a synthetic logic corpus designed to offer broad and systematic coverage of logical knowledge. However, it remains unclear for this important question, RQ2: What style of training data, natural language or formal logical symbols, better facilitates the learning of multi-step logical reasoning through SFT? Addressing this research question is important for understanding how to most effectively instill logical reasoning capabilities in LLMs.
@@
To address RQ1, we propose FineLogic, a diagnostic framework designed to offer a multi-dimensional evaluation of LLMs’ reasoning processes, moving beyond binary final-answer correctness. Rather than creating another leaderboard, FineLogic serves as a tool for LLM practitioners to pinpoint specific weaknesses in reasoning chains, such as flawed logic, redundancy, or non-atomic steps. Our framework evaluates models along three complementary dimensions: (1) Overall benchmark accuracy: This metric captures a model’s ability to perform multi-step logical reasoning and its generalizability across problems from diverse domains. (2) Stepwise Soundness: Inspired by Saparov and He (2022), we assess the quality of each intermediate reasoning step using three criteria—validity (whether the step is logically valid), relevance (whether its conclusion is used in later steps), and atomicity (whether it applies a single, minimal inference rule). These metrics aim to evaluate the model’s ability to generate human-interpretable and logically coherent reasoning chains. (3) Representation-level probing Ye et al. (2024): By applying probing techniques to LLM hidden representations, this evaluation provides insight into whether the model’s understanding of logical structure is merely surface-level or embedded in its internal state.
@@
  • We propose FineLogic, a unified diagnostic framework for assessing LLMs’ logical reasoning. It moves beyond final-answer accuracy to evaluate the quality and coherence of reasoning steps, offering fine-grained insights that can guide targeted model improvements and serve as a scaffold for structured reward design in reinforcement learning.
@@
  • Through systematic analysis of models trained with different supervision styles, we identify key trade-offs between generalization and structural reasoning quality. The findings provide concrete insights into the design and selection of effective training data for post-training.
@@ 2 Related Works
Logical Reasoning Benchmarks. Numerous benchmarks have been proposed to evaluate the logical reasoning abilities of LLMs. Many either mix logical and commonsense reasoning Liu et al. (2023); Luo et al. (2023); Havrilla et al. (2024), making it hard to isolate logical competence, or assess multi-step reasoning using only final-answer accuracy Parmar et al. (2024); Han et al. (2022); Tafjord et al. (2020); Mondorf and Plank (2024). While ProntoQA Saparov and He (2022); Saparov et al. (2023) pioneered stepwise evaluation, its analysis is confined to step correctness on short problems. FineLogic moves beyond this by introducing a more comprehensive suite of step-level diagnostics—including a more practical relevance metric—to rigorously assess complex, long-chain reasoning. Furthermore, we are the first to adapt representation-level probing from mathematics Ye et al. (2024) to the logical domain, introducing novel tasks like Correctness Spanning Steps (CSS) to connect behavioral outputs with internal model states. In contrast, other stepwise frameworks like ROSCOE Golovneva et al. (2022) and RECEVAL Prasad et al. (2023) remain too generic or coarse-grained for the specific demands of formal logic.
Logical Reasoning Enhancement. Several studies have aimed to improve LLMs’ performance on logical reasoning tasks. Some approaches rely on translating inputs into formal logic and using programmable verifiers to solve problems Olausson et al. (2023); Pan et al. (2023); Yang et al. (2023); Ryu et al. (2024), which bypasses the model’s own reasoning process. Others use in-context learning or inference-time strategies to guide output without fundamentally enhancing reasoning ability Creswell et al. (2022); Wang et al. (2025a); Xu et al. (2024); Sun et al. (2023); Toroghi et al. (2024). While a few works have explored fine-tuning or reinforcement learning to strengthen logical reasoning Feng et al. (2023); Morishita et al. (2023, 2024); Xie et al. (2025); Yang et al. (2022); Xie et al. (2024); Zheng et al. (2025), they have not examined which types of supervision are most effective for teaching LLMs to reason. In this work, we focus specifically on this open question.
@@ 3.1 Overall Benchmark Accuracy
Similar to most benchmarks, our overall benchmark accuracy focuses on final-answer correctness. Specifically, we define accuracy (Acc) as:
$\text{Acc}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[\hat{y}_{i}=y_{i}]$ (1)
where $N$ is the number of test problems, $y_{i}$ is the gold label, and $\hat{y}_{i}$ is the model’s prediction for problem $i$. While coarse-grained, it offers a quick and effective way to assess a model’s overall reasoning ability and cross-domain generalization. We evaluate on four challenging multi-step reasoning benchmarks, deliberately selected for their focus on peer-reviewed, multi-hop formal logic. The suite includes widely-used datasets for comparability (FOLIO Han et al. (2022) and ProntoQA Saparov and He (2022)) alongside benchmarks designed to probe out-of-distribution generalization by varying complexity and rule composition (FLD Morishita et al. (2024) and Multi-LogiEval Patel et al. (2024)). This selection offers a comprehensive assessment, in contrast to narrower benchmarks like ProofWriter Tafjord et al. (2020), which we exclude due to its limited scope of logical rules. Further details on dataset statistics and sampling are provided in Table 1 and Appendix A.1.
@@
FLD Morishita et al. (2024) 1100 {T, F, Unknown}
FOLIO Han et al. (2022) 203 {T, F, Unknown}
Multi-Logical Patel et al. (2024) 390 {T, F}
Pronto-QA Saparov and He (2022) 500 {T, F}
@@ 3.2 Stepwise Soundness
To assess these criteria, we extract the premises and conclusion of each step from the model’s final answer. Crucially, our evaluator scores only the text following a designated answer tag (e.g., </think>), ensuring that any preceding self-reflection or speculative reasoning is excluded by design. We use GPT-4.1-mini to evaluate validity and atomicity. Manual verification on 200 annotated steps shows that GPT-4.1-mini achieves over 98% accuracy on both metrics. For relevance, we determine whether the conclusion of step $i$ is referenced in any subsequent step $k>i$.
We then compute the proportion of samples in which all steps are valid, relevant, and atomic, providing a sample-level measure of reasoning integrity. Formally, for each solution $s$ with $K_{s}$ steps, let $v_{s,k},r_{s,k},a_{s,k}\in\{0,1\}$ denote the validity, relevance, and atomicity of step $k$, respectively. We define the sample-level metrics as:
AllValid $=\frac{1}{N}\sum_{s=1}^{N}\left[\prod_{k=1}^{K_{s}}v_{s,k}\right]\quad$ (2)
AllRelevant $=\frac{1}{N}\sum_{s=1}^{N}\left[\prod_{k=1}^{K_{s}}r_{s,k}\right]\quad$ (3)
AllAtomic $=\frac{1}{N}\sum_{s=1}^{N}\left[\prod_{k=1}^{K_{s}}a_{s,k}\right]$ (4)
Full prompt templates are provided in Figures 13 and 14.
@@ 3.3 Representation-level Probing
Correctness Spanning Steps (CSS): Identifies the earliest step after which the model consistently predicts the correct label. The spanning length is the number of remaining steps from that point to the end. Higher accuracy indicates earlier internalization of the correct answer. Formally, for each problem $s$, let $\tau_{s}$ be the first step from which the probe consistently predicts the correct label and $K_{s}$ be the total number of steps. The CSS score is then:
$\text{CSS}=\frac{1}{N}\sum_{s=1}^{N}(K_{s}-\tau_{s}).$ (5)
@@
For both RFI and NSD, performance is measured using balanced accuracy to account for the construction of positive and negative instances. The score is calculated as:
$\text{Score}=\frac{1}{2}(\text{TPR}+\text{TNR}),$
where TPR is the True Positive Rate and TNR is the True Negative Rate.
Our evaluation builds on two prior lines of work—stepwise reasoning evaluation Saparov and He (2022) and representation-level probing Ye et al. (2024)—but introduces key extensions tailored to logical reasoning.
@@ 5.1 Experimental Setup
Our experiments involve four base models for fine-tuning—LLaMA-3.1-8B-Instruct, Qwen-2.5-7B-Instruct, Qwen-2.5-Math-7B-Instruct, and Qwen-3-8B—and two powerful zero-shot baselines, GPT-4o and DeepSeek R1. The four base models are fine-tuned for 3 epochs at a $1\times 10^{-6}$ learning rate, with their original versions also serving as baselines. Representation-level probing is limited to the LLaMA and Qwen models due to computational constraints, and an explicit step-by-step format is enforced for all stepwise evaluations.
@@
  • Few-Shot Learning Brown et al. (2020)
  • LOGIPT (Creswell et al., 2022)
  • Selection-Inference (Creswell et al., 2022)
  • SymbCoT Xu et al. (2024)
  • LogicLM Pan et al. (2023)
@@
[…]
@@ 5.2 Results
We conducted experiments for analyzing the performance of four models combined with various prompting and fine-tuning settings under the FineLogic Evaluation Framework. Due to space constraints, the results for Qwen-2.5-Math-7B-Instruct and Qwen-3-8B are presented in Appendix C.
@@ 5.2.1 Results on Overall Benchmark Accuracy
As shown in Table 1, we report the overall benchmark accuracy across four datasets, as well as the step-wise accuracy on the FLD benchmark, stratified by reasoning depth (Figure 3). Our analysis yields several key observations:
@@ 5.2.2 Results on Stepwise Soundness
The All Valid metric measures the proportion of samples where every step is logically valid, a stringent indicator of formal reasoning. We observe that models fine-tuned with SFT-NL and SFT-Symb-Struct achieve high All Valid scores. Llama with SFT-NL reaches 40.9%, substantially outperforming strong baselines like GPT-4o and DeepSeek-R1. The low score of DeepSeek-R1, despite its high accuracy, highlights a key insight from FineLogic. Having been trained solely on final-answer correctness, it often compresses multiple reasoning steps and omits necessary premises for intermediate conclusions. Since guessing the final label is easier than ensuring every inference is sound, accuracy alone can be misleading. "All Valid" enforces this stricter requirement, revealing a critical gap in the model’s reasoning fidelity.
@@
5.2.3 Results on Representation-level Probing
Table 4 presents results from our probing analysis, which assesses whether models internally acquire key reasoning abilities. Our findings reveal several key takeaways:
Correctness Spanning Steps (CSS) shows a clear ceiling for 7/8B models under SFT. Across all SFT variants and base models, the CSS score fluctuates by less than 1%. This flat trend suggests that supervised fine-tuning alone offers little headroom for improving how early a model internally converges on the correct answer at this scale. Enhancing this "foresight" capability may require alternative methods like reinforcement learning with step-level rewards or architectural changes.
Targeted supervision can unlock specific internal skills, as shown by Symb-Direct on RFI. The Redundant Fact Identification (RFI) metric benefits most from the SFT-Symb-Direct setting. This training data uses minimal symbolization but retains redundant facts, forcing the model to learn to distinguish necessary from unnecessary premises. This direct alignment between the training task and the probing metric demonstrates that when a specific skill like redundancy filtering is desired, targeted supervision that mirrors the probe is highly effective.
Next-Step Derivability (NSD) remains a challenge. Performance on the Next-Step Derivability (NSD) task shows no consistent improvement across SFT variants. This indicates that vanilla SFT does not effectively enhance the model’s internal representation of what logical steps are derivable next. This highlights a remaining gap in teaching explicit, forward-looking inference, suggesting promising future directions such as combining the NSD probe with self-supervised next-step prediction or curriculum-based RL.
6 Future Work: Guiding Faithful Reasoning with Reinforcement Learning
While supervised fine-tuning improves reasoning, reinforcement learning (RL) offers a path to refining it further. However, a major challenge in applying RL to logical reasoning is the reliance on sparse, binary rewards (i.e., whether the final answer is correct). This can lead to "reward hacking," where models learn heuristics to guess the right answer without developing a sound reasoning process.
The FineLogic framework offers a direct solution by providing dense, multi-faceted reward signals that can guide the quality of the reasoning process itself. Because our metrics are reference-free and deterministically scored, they translate cleanly into a multi-objective reward function.
A Multi-Objective Reward Framework.
The FineLogic metrics can be directly integrated into a granular, multi-objective reward function for RL. The stepwise soundness metrics (validity, relevance, atomicity) can be used to directly reward logical integrity, promoting reasoning that is rigorous, concise, and transparent. Concurrently, the representation-level metric CSS can serve as an online reward for cognitive efficiency, encouraging the model to converge on the correct answer earlier in its internal process. These components can be combined in a weighted sum:
$R_{\text{total}}=$ $R_{\text{acc}}+w_{v}R_{\text{valid}}+w_{r}R_{\text{relevant}}$
$+w_{a}R_{\text{atomic}}+w_{c}R_{\text{css}}$ (6)
In this framework, practitioners can tune the weights ($w_{v},w_{r},\dots$) to prioritize desired reasoning qualities like rigor or conciseness, offering a practical roadmap for training more faithful and interpretable models.